find the slope of the graph??
(1988,1713), (2003,5671)
: find the slope of the graph??
(1988,1713), (2003,5671)

Answer by midwood_trail(310)

Find the slope of the graph?
(1988,1713), (2003,5671)
I will use the letter S for slope.
To find the slope S, use this formula:
S = (d - b)/(c - a)
In the given two points, a = 1988, b = 1713, c = 2003, d = 5671.
Let's plug and chug.
S = (5671 - 1713)/(2003 - 1988)
S = 3958/15
Did you follow?